A 4-year maturity treasury note has a yield to maturity of 4%. Its par value is $1,000, it makes semi-annual payments (twice per year), and its coupon rate is 3%. What is this bond’s current price? The first coupon arrives in 6 months.
Answer #1

Price of a bond is mathematically represented as:

where P is price of bond, with periodic coupon C, n periods to maturity, periodic YTM i and face value M.

For this question,

n = 4 * 2 = 8 semi-annual periods,

C = 3% * $1000/2 = $15 (semi-annually),

i = 4%/2 = 2% (semi-annually)

M = $1000

P = $109.88 + $853.49

P = $963.37